# Towards Adversarially Robust Dataset Distillation by Curvature Regularization

### v1

@@ 6.2 Computational Overhead
6.3 Transferability
Our investigation focuses on studying the effectiveness of the curvature regularizer within the SRe2L framework. Theoretically, this method can be extended to a broad spectrum of dataset distillation methods. GUARD’s application is feasible for any distillation approach that utilizes a model trained on the original dataset as a comparison target during the distillation phase — a strategy commonly seen across many dataset distillation techniques as noted in Section 2. This criterion is met by the majority of dataset distillation paradigms, with the exception of those following the distribution matching approach, which may not consistently employ a comparison model (Sachdeva and McAuley, 2023). This observation suggests GUARD’s potential compatibility with a wide array of dataset distillation strategies.
To demonstrate this, we explored two additional implementation of GUARD using DC (Zhao et al., 2021) and CDA (Yin and Shen, 2023) as baseline distillation methods. DC represents an earlier, simpler approach that leverages gradient matching for distillation purposes, whereas CDA is a more recent distillation technique, specifically designed for very large datasets. As shown in Table 5, GUARD consistently improves both clean accuracy and robustness across various dataset distillation methods.
Table 5: Direct comparison of the original DC, SRe2L, and CDA methods with the addition of GUARD regularizer on CIFAR10. The best results among each pair of compared methods are highlighted in bold.
@@ 6.3 Transferability
7 Conclusion
Our work focuses on a novel perspective on dataset distillation by emphasizing its robustness characteristics. Upon reaching the theoretical conclusion that the adversarial loss of distilled datasets is bounded by the curvature, we proposed GUARD, a method that can be integrated into many dataset distillation methods to provide robustness against diverse types of attacks and potentially improve clean accuracy. Our theory also provided the insight that the optimization of robustness with respect to distilled and real datasets is differentiated only by a constant term. This conclusion opens up various potentials for subsequent research in the field. We hope that our work contributes to the development of dataset distillation techniques that are not only efficient but also robust, and will inspire and support further research and development in this area.
@@ Appendix A Proof of Proposition 1
~$\begin{split}\tilde{\ell}_{\rho}^{adv}(\mathbf{x})&=\max_{\|\mathbf{v}\|\leq\rho}\tilde{\ell}(\mathbf{x+v})=\max_{\|\mathbf{v}\|\leq\rho}\ell(\mathbf{x})+{\nabla\ell(\mathbf{x})}^{\top}\mathbf{v}+\frac{1}{2}\mathbf{v^{\top}Hv}\\ &\leq\max_{\|\mathbf{v}\|\leq\rho}\ell(\mathbf{x})+\|{\nabla\ell(\mathbf{x})}\|\|\mathbf{v}\|+\frac{1}{2}\lambda_{1}(\mathbf{x}){\|\mathbf{v}\|}^{2}\\ &=\ell(\mathbf{x})+\|{\nabla\ell(\mathbf{x})}\|\rho+\frac{1}{2}\lambda_{1}(\mathbf{x})\rho^{2},\end{split}$ (12)
@@
$\mathop{\mathbb{E}}_{\mathbf{x}\sim D_{c}}\tilde{\ell}_{\rho}^{adv}(\mathbf{x})\leq\mathop{\mathbb{E}}_{\mathbf{x}\sim D_{c}}\ell(\mathbf{x})+\rho\mathop{\mathbb{E}}_{\mathbf{x}\sim D_{c}}\|{\nabla\ell(\mathbf{x})}\|+\frac{1}{2}\rho^{2}\mathop{\mathbb{E}}_{\mathbf{x}\sim D_{c}}\lambda_{1}(\mathbf{x}).$ (13)
@@
~$\tilde{\ell}_{\rho}^{adv}(\mathbf{x}^{\prime})\leq\mathop{\mathbb{E}}_{\mathbf{x}\sim D_{c}}\ell(\mathbf{x})+\rho\mathop{\mathbb{E}}_{\mathbf{x}\sim D_{c}}\|\nabla\ell(\mathbf{x})\|+\frac{1}{2}\rho^{2}\mathop{\mathbb{E}}_{\mathbf{x}\sim D_{c}}\lambda_{1}(\mathbf{x})+L\sigma.$ (18)
@@ Discussion
The inequality in line (2) is an equality if and only if the direction of the gradient is the same as the direction of $\lambda_{1}$. Previous work has empirically shown that the two directions have a large cosine similarity in the input space of neural networks. Our assumption about the Lipschitz continuity of $\mathcal{L}_{\rho}^{adv}(\cdot)$ is reasonable, as recent work has shown improved estimation of the Lipschitz constant of neural networks in a wide range of settings (Khromov and Singh, 2023). Although our assumptions about the convexity of ~$\tilde{\ell}(\mathbf{x})$ and the linearity of $h(\cdot)$ is relatively strong, it still reflects important aspects of reality, as our experiment in Table 2 has shown that reducing the curvature term in r.h.s of  Eq. 18 effectively improves the robustness of models trained on distilled data. Moreover, in Fig. 2 we plot the distribution of eigenvalues of the real data samples on the loss landscape of a model trained on standard distilled data and a model trained on robust distilled data from our GUARD method, respectively. GUARD corresponds to a flatter curve of eigenvalue distribution, indicating that the loss landscape becomes more linear after our regularization.
@@
Appendix B Effect on GUARD on Images
@@
Appendix C GUARD and Optimization-based Distillation Methods
In Section 4.4, we explained how GUARD can be easily intergrated into SRe2L by incorporating the regularizer into the model’s training loss during the squeeze (pre-training) step. We later demonstrated that GUARD could also be integrated into other distillation methods, such as DC (Zhao et al., 2021). However, unlike SRe2L, DC lacks a pre-training phase; instead, the model’s training and distillation occur simultaneously, making the integration less straightforward.
Therefore, we present the GUARD algorithm using DC as the baseline method in Algorithm 1. For each outer iteration $k$, we sample a new initial weight from some random distribution of weights to ensure the synthetic dataset can generalize well to a range of weight initializations. After, we iteratively sample a minibatch pair from the real dataset and the synthetic dataset and compute the loss over them on a neural network with the weights $\boldsymbol{\theta}_{t}$. We compute the regularized loss on real data through Eq. 11. Finally, we compute the gradient of the losses w.r.t. $\boldsymbol{\theta}$, and update the synthetic dataset through stochastic gradient descent on the distance between the gradients. At the end of each inner iteration $t$, we update the weights $\boldsymbol{\theta}_{t+1}$ using the updated synthetic dataset.
Input: $\mathcal{T}$: Training set; $\mathcal{S}$: initial synthetic dataset with $C$ classes; $p(\theta_{0})$: initial weights distribution; $\phi_{\theta}$: neural network; $K$: number of outer-loop steps; $T$: number of inner-loop steps; $\varsigma_{\theta}$: number of steps for updating weights; $\varsigma_{S}$: number of steps for updating synthetic samples; $\eta_{\theta}$: learning rate for updating weights; $\eta_{S}$: learning rate for updating synthetic samples; $D$: gradient distance function; $h$: discretization step; $\lambda$: strength of regularization for each outer training step $k=1$ to $K$ do
@@ Appendix C GUARD and Optimization-based Distillation Methods
Algorithm 1 DC-based GUARD

### v4

@@ Discussion
Transferability
Our investigation focuses on studying the effectiveness of the curvature regularizer within the SRe2L framework. Theoretically, this method can be extended to a broad spectrum of dataset distillation methods. GUARD’s application is feasible for any distillation approach that utilizes a model trained on the original dataset as a comparison target during the distillation phase — a strategy commonly seen across many dataset distillation techniques as noted in the Related Works section. This criterion is met by the majority of dataset distillation approaches, with the exception of those following the distribution matching approach, which may not consistently employ a comparison model (Sachdeva and McAuley 2023). This observation suggests GUARD’s potential compatibility with a wide array of dataset distillation strategies. To demonstrate this, we explored two additional implementation of GUARD using DC (Zhao, Mopuri, and Bilen 2021) and CDA (Yin and Shen 2023) as baseline distillation methods. DC represents an earlier, simpler approach that leverages gradient matching for distillation purposes, whereas CDA is a more recent distillation technique, specifically designed for very large datasets. As shown in Table 5, GUARD consistently improves both clean accuracy and robustness across various dataset distillation methods.
Table 5: Direct comparison of the original DC, SRe2L, and CDA methods with the addition of GUARD regularizer (marked by ) on CIFAR10. The best results among each pair of compared methods are highlighted in bold.
@@
Conclusions
Our work focuses on a novel perspective on dataset distillation by emphasizing its adversarial robustness characteristics. Upon reaching the theoretical conclusion that the adversarial loss of distilled datasets is bounded by the curvature, we proposed GUARD, a method that can be integrated into many dataset distillation methods to provide robustness against diverse types of attacks and potentially improve clean accuracy. Our theory also provided the insight that the optimization of robustness with respect to distilled and real datasets is differentiated only by a constant term, which may open up potentials for subsequent research in the field. Future work could explore the integration of robustness into more dataset distillation approaches as well as out-of-distribution settings. We hope our work contributes to the development of DD methods that are not only efficient but also robust, and will inspire further research in this area.
@@ Conclusions
~$\begin{split}\tilde{\ell}_{\rho}^{adv}(\mathbf{x})&=\max_{\|\mathbf{v}\|\leq\rho}\tilde{\ell}(\mathbf{x+v})\\ &=\max_{\|\mathbf{v}\|\leq\rho}\ell(\mathbf{x})+{\nabla\ell(\mathbf{x})}^{\top}\mathbf{v}+\frac{1}{2}\mathbf{v^{\top}Hv}\\ &\leq\max_{\|\mathbf{v}\|\leq\rho}\ell(\mathbf{x})+\|{\nabla\ell(\mathbf{x})}\|\|\mathbf{v}\|+\frac{1}{2}\lambda_{1}(\mathbf{x}){\|\mathbf{v}\|}^{2}\\ &=\ell(\mathbf{x})+\|{\nabla\ell(\mathbf{x})}\|\rho+\frac{1}{2}\lambda_{1}(\mathbf{x})\rho^{2},\end{split}$ (16)
@@
$\begin{split}\mathop{\mathbb{E}}_{\mathbf{x}\sim D_{c}}\tilde{\ell}_{\rho}^{adv}(\mathbf{x})\leq\mathop{\mathbb{E}}_{\mathbf{x}\sim D_{c}}\ell(\mathbf{x})+\rho\mathop{\mathbb{E}}_{\mathbf{x}\sim D_{c}}\|{\nabla\ell(\mathbf{x})}\|\\ +\frac{1}{2}\rho^{2}\mathop{\mathbb{E}}_{\mathbf{x}\sim D_{c}}\lambda_{1}(\mathbf{x}).\end{split}$ (17)
@@
~$\begin{split}\tilde{\ell}_{\rho}^{adv}(\mathbf{x}^{\prime})\leq\mathop{\mathbb{E}}_{\mathbf{x}\sim D_{c}}\ell(\mathbf{x})+\rho\mathop{\mathbb{E}}_{\mathbf{x}\sim D_{c}}\|\nabla\ell(\mathbf{x})\|\\ +\frac{1}{2}\rho^{2}\mathop{\mathbb{E}}_{\mathbf{x}\sim D_{c}}\lambda_{1}(\mathbf{x})+L\sigma.\end{split}$ (22)
@@ Discussion
The inequality in Eq. 16 is an equality if and only if the direction of the gradient is the same as the direction of $\lambda_{1}$. Previous work has empirically shown that the two directions have a large cosine similarity in the input space of neural networks. Our assumption about the Lipschitz continuity of $\mathcal{L}_{\rho}^{adv}(\cdot)$ is reasonable, as recent work has shown improved estimation of the Lipschitz constant of neural networks in a wide range of settings (Khromov and Singh 2023). Although our assumptions about the convexity of ~$\tilde{\ell}(\mathbf{x})$ and the linearity of $h(\cdot)$ is relatively strong, it still reflects important aspects of reality, as our experiment in Table 2 has shown that reducing the curvature term in r.h.s of Eq. 22 effectively improves the robustness of models trained on distilled data. Moreover, in Fig. 2 we plot the distribution of eigenvalues of the real data samples on the loss landscape of a model trained on standard distilled data and a model trained on robust distilled data from our GUARD method, respectively. GUARD corresponds to a flatter curve of eigenvalue distribution, indicating that the loss landscape becomes more linear after our regularization.
@@
Additional Results
In this section of the appendix, we provide supplementary results from our experiments. Table 6 presents a detailed comparison of the effects of various adversarial attacks on GUARD, SRe2L, MTT, and TESLA, which were excluded from the main paper due to space constraints. The results further highlight GUARD’s improved adversarial robustness, mai=ntaining a positive trend of being much better than other methods.
Furthermore, Table 7 showcases additional comparisons illustrating the computational efficiency of GUARD. The results demonstrate that GUARD consistently achieves significantly faster runtimes per iteration compared to adversarial training.
Table 6: Evaluation of different dataset distillation methods under adversarial attacks on ImageNette and TinyImageNet, under the 1 IPC setting. The best results among all methods are highlighted in bold, second best are underlined.
Dataset IPC Attack Methods
GUARD SRe2L MTT TESLA
Imagenette 1 PGD100 16.22 12.05 18.60 24.68
Square 26.74 18.62 1.40 22.21
AutoAttack 15.81 12.12 1.40 22.16
CW 29.14 20.38 1.40 22.26
MIM 16.32 12.05 18.60 24.68
TinyImagenet 1 PGD100 1.48 0.83 3.26 3.64
Square 3.26 2.44 1.76 2.83
AutoAttack 0.74 0.66 1.76 2.78
CW 1.54 1.18 1.76 2.82
MIM 1.49 0.88 3.28 3.64
Table 7: Relative slowdown introduced by adding GUARD versus adversarial training to SRe2L in terms of average runtime per iteration, tested on three variations of the ImageNette dataset with image sizes of 160x160px, 320x320px, and the original size, using two different graphics cards.
Method A100 RTX4090
160px 320px Original 160px 320px Original
GUARD 4.63x 3.96x 4.54x 2.37x 3.26x 2.81x
Adv. Training 49.83x 49.96x 53.40x 35.14x 53.22x 41.23x
Effect of GUARD on Images
@@
Algorithm of GUARD with Optimization-based Distillation Methods
In our main paper, we explained how GUARD can be easily intergrated into SRe2L by incorporating the regularizer into the model’s training loss during the squeeze (pre-training) step. We later demonstrated that GUARD could also be integrated into other distillation methods, such as DC (Zhao, Mopuri, and Bilen 2021). However, unlike SRe2L, DC lacks a pre-training phase; instead, the model’s training and distillation occur simultaneously, making the integration less straightforward.
Therefore, we present the GUARD algorithm using DC as the baseline method in Alg. 1. For each outer iteration $k$, we sample a new initial weight from some random distribution of weights to ensure the synthetic dataset can generalize well to a range of weight initializations. After, we iteratively sample a minibatch pair from the real dataset and the synthetic dataset and compute the loss over them on a neural network with the weights $\boldsymbol{\theta}_{t}$. We compute the regularized loss on real data through Eq. 15. Finally, we compute the gradient of the losses w.r.t. $\boldsymbol{\theta}$, and update the synthetic dataset through stochastic gradient descent on the distance between the gradients. At the end of each inner iteration $t$, we update the weights $\boldsymbol{\theta}_{t+1}$ using the updated synthetic dataset.
Input: $\mathcal{T}$: Training set; $\mathcal{S}$: initial synthetic dataset with $C$ classes; $p(\theta_{0})$: initial weights distribution; $\phi_{\theta}$: neural network; $K$: number of outer-loop steps; $T$: number of inner-loop steps; $\varsigma_{\theta}$: number of steps for updating weights; $\varsigma_{S}$: number of steps for updating synthetic samples; $\eta_{\theta}$: learning rate for updating weights; $\eta_{S}$: learning rate for updating synthetic samples; $D$: gradient distance function; $h$: discretization step; $\lambda$: strength of regularization
for each outer training step $k=1$ to $K$ do
@@
Algorithm 1 Algorithm of GUARD (based on DC)